OpenStudy (anonymous):

Translate this to a mathematical equation and solve. Please show all of your work. A bike road race starts at an elevation of 870 feet and passes through five stages where the elevation changes by -684 feet, 17 feet, -236 feet, -440 feet, and 286 feet. At what elevation does the race end?

OpenStudy (radar):

870-684+17-236-440+286=??????

OpenStudy (radar):

870+17+286-(684+236+440)= 1173-1360=-187 The racers are at 870-187 or at 683 elevation

OpenStudy (anonymous):

you already started at 870 in your equation so you added an extra 870 feet I think. The final elevation is -187 feet.

OpenStudy (anonymous):

Yes...that's what I got.

OpenStudy (anonymous):

Thanks

OpenStudy (radar):

Agree, I goofed. I guess they are below sea level lol